A jet airliner moving initially at 2.85 x 102 mi/h
Answer #1

Break the wind down into components (north and east) and add the wind speed to the airliner's speed.

285 + 100cos(33.5°) = 368.39 mph E
0 + 100sin(33.5°) = 55.19 mph N

Then use the pythagorean theorem to find the magnitude, and tangent to find the direction.

√[368.39² + 55.19²] = 372.501 mph
tanθ = 55.19 / 368.39 = ~ 0.1498

Calculations maybe wrong, logic is correct.

Hope this helps !! :)